The wheel starts spinning evenly and at n = 10 turns w = 20rad / s
the angels speeded up. Find the angular acceleration ε of a circle.
1
Expert's answer
2020-03-02T10:21:45-0500
"\omega=\alpha t"

"\theta=2\pi n=0.5\alpha t^2"

"2\pi n=0.5\alpha \left(\frac{\omega}{\alpha}\right)^2"

"4\pi n=\left(\frac{\omega^2}{\alpha}\right)"

"4\pi (10)=\left(\frac{20^2}{\alpha}\right)"

"\alpha=3.18\frac{rad}{s^2}"